OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 15 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 725 miles, for a total gas consumption of 35 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

20*25=500 15*15=225 500+225=725

OpenStudy (anonymous):

20 gallons in the more efficient car, and 15 gallons in the less efficient car.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

Your welcome.